# Dialogue Without Limits: Constant-Sized KV Caches for Extended Responses in LLMs

### v1

@@ 5.3.2 KV cache Sizes
6 Ablation studies
The design of MorphKV consists of three key design parameters, namely, fusion function used to create the attention profile, recent window size, and total KV cache budget. These hyper-parameters collectively influence the final benchmark performance with MorphKV. In this section, we analyze the impact of each hyper-parameter by experimenting with fusion functions, testing various window sizes, and substantially compressing the KV cache budget for MorphKV, while drawing comparisons with existing methods.
6.1 Impact of Fusion Function: sum() Versus max()
6.1.1 On Long-Context Tasks
6.1.2 On Long-Response Tasks
@@
6.2 Impact of Window Size on Long-Response Tasks
We evaluate the impact of window size on the overall performance of MorphKV. Intuitively, recent-window tokens are used to select which tokens to retain from the distant past. Hence, a larger window allows capturing of more diverse information from the past.
@@ 6.2 Impact of Window Size on Long-Response Tasks
Table 6: LongGenBench: Sensitivity of evaluation metrics with window size across different models (CR: Completion Rate)
As shown in Table 6, changing the window size significantly impacts evaluation metrics for both the Llama and Mistral models. This is due to the fact that a very small window does not suffice for capturing extensive amounts of distant information present in a typical LongGenBench prompt (such as specific details about different floors in a building, specific Menu items etc.), because very small windows tend to capture local context, while failing to capture more distant context. Therefore, a larger window is more effective since it allows the model to retain diverse pieces of information even at very large response sizes. For instance, a window of size 200 lets Llama3.1 recall accurate instructions regarding the 96th floor of a building, in spite of already generating extensive descriptions of the previous 95 floors. On the other hand, with a window size of 32 tokens, the model struggles to maintain consistency with the input request, and generates generic responses after certain number of floors, thereby losing on accuracy.
6.3 Robustness Against KV Cache Compression
@@
Figure 10: LongBench: Average scores for (a) Llama3.1-8B and (b) Mistral-7B models, comparing MorphKV and SnapKV under different levels of KV cache compression. SnapKV degrades much more sharply with decreasing KV cache budget, while MorphKV offers stable performance even at a low KV cache budget.
Figure 10 shows the mean benchmark score (we refer to Appendex A.1 for individual benchmark scores) for both models under varying compression scenarios. At very low KV cache budgets, we observe a drop of more than $50\%$ on average between MorphKV and SnapKV, for benchmarks like NarrativeQA, this difference reaches upto $88\%$. This disparity indicates that MorphKV is significantly more effective at retaining crucial context information compared to SnapKV under tight memory constraints. Even with larger budgets, MorphKV consistently outperforms SnapKV, demonstrating the robustness and reliability of its design.
@@
7 Conclusion
The growing memory footprint of KV caches in LLMs poses a critical bottleneck for long-context and long-response tasks. In this paper, we propose MorphKV that addresses this challenge by introducing a dynamic, correlation-aware token selection mechanism that maintains a constant-sized KV cache while preserving contextual coherence. MorphKV leverages attention profiles of recent tokens to identify and retain only the most relevant distant tokens. Our studies on long-response tasks show 52.9% memory savings and 18.2% higher accuracy on average compared to state-of-the-art prior works. Our experiments demonstrate that MorphKV scales efficiently with response length, degrading only 10% in performance even as outputs grow to 12K tokens, compared to 15–18% degradation for state-of-the-art prior works, SnapKV and H2O. Furthermore, MorphKV’s compatibility with Grouped Query Attention enables 4$\times$ greater memory efficiency than Multi Headed Attention-based approaches, making it practical for real-world deployment in chatbots, interactive assistants, and content generation systems. These advances position MorphKV as a practical inference-time solution for LLMs, balancing both accuracy and memory constraints without sacrificing the ability to capture long-range dependencies.
Acknowledgments
The authors acknowledge the Texas Advanced Computing Center (TACC) and the Center for Generative AI at the University of Texas at Austin for providing computational resources that helped develop the research results reported in this paper. Poulami Das acknowledges the generous support through the AMD endowment at the University of Texas at Austin. Prashant J. Nair is supported by Intel Transformation Server Architecture (TSA) and the Natural Sciences and Engineering Research Council of Canada (NSERC) [funding reference number RGPIN-2019-05059] Grants. We acknowledge Won Joon Yun for editorial feedback.
Appendix A Appendix
A.1 LongBench: Comparison of MorphKV and SnapKV under different KV cache budget constraints
We stress-test both MorphKV and SnapKV for robustness against KV cache compression, by varying the compression rates on LongBench. Figure 11 shows per dataset trends, we note that MorphKV consistently provides robust performance even at extreme levels of KV cache compression, while SnapKV’s performance degrades drastically at very low KV cache budget.
Figure 11: Comparison of MorphKV versus SnapKV for Llama3.1-8B, and Mistralv0.2-7B. (a)-(c) show results for Llama3.1-8B on MultifieldQA-en, Musique, and NarrativeQA respectively, while (d)-(f) corresponds to Mistralv0.2-7B evaluation for NarrativeQA, PassageRetrieval-en, and 2wikimqa respectively. MorphKV consistently beats SnapKV across varying KV cache budget sizes.
A.2 LongWriter: LLM Judge Scores for Different Models and Metrics
Table 7 provides detailed LLM Judge Scores across various criteria. All configurations use a KV cache capacity of 600 tokens, with a recent window size of 30 tokens for both SnapKV and MorphKV. MorphKV consistently outperforms H2O on Relevance, highlighting its effectiveness in retaining imporatant tokens. Qwen2.5 differs from other models as it employs GQA with 7×7$\times$7 × less heads than the default MHA configuration, potentially introducing partial information loss when MorphKV operates under fewer heads, explaining why MorphKV performs compared to SnapKV for this model.

### v2

@@ A.1.1 On Long-Context Tasks
A.1.2 On Long-Response Tasks
@@
A.2 Impact of Window Size on Long-Response Tasks
We evaluate the impact of window size on MorphKV’s performance. Intuitively, recent-window tokens determine which tokens to retain from the distant past. Thus, a larger window allows capturing of more diverse information from the past.
@@ A.2 Impact of Window Size on Long-Response Tasks
As shown in Table 8, changing the window size significantly impacts evaluation metrics for all Llama, Mistral, and Qwen models. This is due to the fact that a very small window does not suffice for capturing extensive amounts of distant information present in a typical LongGenBench prompt (such as specific details about different floors in a building, specific Menu items etc.), because very small windows tend to capture local context, while failing to capture more distant context. Therefore, a larger window is more effective since it allows the model to retain diverse pieces of information even at very large response sizes. For instance, a window of size 200 lets Llama3.1 recall accurate instructions regarding the 96th floor of a building, in spite of already generating extensive descriptions of the previous 95 floors. On the other hand, with a window size of 32 tokens, the model struggles to maintain consistency with the input request, and generates generic responses after certain number of floors, thereby losing on accuracy.
@@
Table 8: LongGenBench: Sensitivity of evaluation metrics with window size across different models (CR: Completion Rate). A small window is insufficient for capturing large-amounts of distant information in long-response tasks, leading to a degradation in model accuracy for all models. However, Completion rate remains comparable across window sizes as it only measures the model’s ability to complete the response as expected, without considering the relevance or correctness of the generated output.
A.3 Robustness Against KV Cache Compression
@@
Figure 10 shows the mean and individual benchmark scores for both models under varying compression scenarios. At very low KV cache budgets, we observe a drop of more than $50\%$ on average between MorphKV and SnapKV, for benchmarks like NarrativeQA, this difference reaches upto $88\%$. This disparity indicates that MorphKV is significantly more effective at retaining crucial context information compared to SnapKV under tight memory constraints. Even with larger budgets, MorphKV consistently outperforms SnapKV, demonstrating the robustness and reliability of its design.
@@
Figure 10: Comparison of MorphKV versus SnapKV for Llama3.1-8B, and Mistralv0.2-7B. (a)-(d) show results for Llama3.1-8B on Average, MultifieldQA-en, Musique, and NarrativeQA respectively, while (e)-(h) show results for Mistralv0.2-7B on Average, NarrativeQA, PassageRetrieval-en, and 2wikimqa respectively. MorphKV consistently beats SnapKV across varying KV cache budgets.
A.4 LongWriter: LLM Judge Scores for Different Models and Metrics
Table 9 provides detailed LLM Judge Scores across various criteria. All configurations use a KV cache capacity of 600 tokens, with a recent window size of 30 tokens for both SnapKV and MorphKV. MorphKV consistently outperforms H2O on Relevance, highlighting its effectiveness in retaining imporatant tokens. Qwen2.5 differs from other models as it employs GQA with 7×7$\times$7 × less heads than the default MHA configuration, potentially introducing partial information loss when MorphKV operates under fewer heads, explaining why MorphKV performs poorer compared to SnapKV for this model.
@@ A.4 LongWriter: LLM Judge Scores for Different Models and Metrics
A.5 Coarse-Grained Token-Eviction
MorphKV performs a dynamic token eviction at every generation step to maintain a constant-sized KV cache. However, this fine-grained approach incurs runtime overhead as the algorithm must run at every timestep, necessitating memory accesses to load the attention profile. While MorphKV reduces this overhead by employing dedicated CUDA streams to prefetch the attention profile, further reductions could be achieved by invoking the algorithm at certain time intervals. Table 10 presents the performance impact when the token-selection policy is applied at every 8th generation-step (eg., 8, 16, 32 etc.). Adopting a coarser token-eviction results in runtime improvement, albeit at reduced accuracy. This trade-off suggests that the balance between latency, and accuracy can be dynamically tuned to suit task-specific requirements at runtime.
xMorphKVx Completion Rate Avg. Accuracy Runtime
default 72.1% 47.0% 1$\times$
coarse-grained eviction (step=8) 71.9% 46.1% 0.82$\times$
Table 10: LongGenBench: Performance of MorphKV with different token eviction policies on Mistral-7B. Coarse-grained selection incurs a slight degradation in long-response generation but yields better runtime compared to default (per-timestep) eviction.
A.6 Selective Layer Retention with MorphKV
MorphKV only preserves those distant tokens which are heavily attended by the recent window tokens. However, KV cache can also be optimized across attention layers. Prior works (Cai et al., 2024; Wang et al., 2024) have shown that early layers capture more critical information than later ones. Building on this insight, we perform a preliminary experiment where we disable MorphKV for the first three layers i.e., the first three layers are retained for all tokens, to emulate a layer-sensitive strategy. Table 11 shows that this configuration outperforms the default MorphKV configuration, although with an additional 10% memory footprint.
While integrating MorphKV with orthogonal techniques which optimize the KV cache across attention layers (Cai et al., 2024; Wang et al., 2024), and attention heads (Fu et al., 2024; Feng et al., 2024) can further improve performance, such integration necessitates a comprehensive understanding of their trade-offs, and careful parameter tuning. We defer this exploration to future work.
xMorphKV)x Completion Rate Avg. Accuracy
default 70.51% 44.2%
layer-aware optimization 71.96% 46.1%
Table 11: LongGenBench: Performance of MorphKV improves with selection (first three layers) layer retention on Mistral-7B, although at an additional 10% memory overhead. Default configuration of MorphKV has a capacity of 2000 tokens.